OpenStudy (anonymous):

On the graph of the equation 2x − y = 10, what is the value of the y-intercept? -10 -5 5 10

OpenStudy (aum):

Rearrange the given equation so it looks like the slope-intercept form, y = mx + b. Then 'b' will be the y-intercept.

OpenStudy (anonymous):

to get the y intercept you need to know the formula y=mx+b b=the y intercept so first get the y to one side so subtract 2x and move it to the other side so it would look like -y=2x+10 then to get rid of the -y multiply each side by -1 so then it would look like y=-2x-10 then the -10= where the b is so 10 is your y intercept ~ i hope i helped!!!

OpenStudy (anonymous):

would it be -10 or 10?